Let random variable X follow a triangular distribution with lower bound a = 0, most likely value m = 9.2, and upper bound b = 16.

Find P(X > 3.0).

Round your answer to three decimal places of accuracy.
Answer #1

Solution :

Given that,

a = 0

b = 16

P(x > c) = (b - c) / (b - a)

P(x >3.0 ) = (16 - 3.0) / (16 - 0 )

= 13 / 16

= 0.812